OpenStudy (anonymous):

There were 150 freshman and sophomores at the high school musical Reese and her friends attended. Freshman tickets were $2.00 each and sophomore tickets were $3.00 each. The receipts for freshman and sophomores tickets totaled $410. How many sophomores bought tickets? 40 90 110 120

OpenStudy (anonymous):

you have to set up two equations. 150 = x+y and 410 = 2x +3y

OpenStudy (anonymous):

x being number of freshman and y number of sophomores

OpenStudy (anonymous):

and then solve for y

OpenStudy (anonymous):

you have to multiply the (150=x+y) by -2 and then add the two equations together

OpenStudy (anonymous):

its 110